Find the mass of water in 300. grams of a 2.80m KCl solution. The answer is 248 g H2O but I'm not sure how to work out the problem.


1
Expert's answer
2021-09-06T02:35:35-0400

2.80m solution means that 2.80 moles of KCl is dissolved per 1 kg of solvent.
This means that "2.80\ moles\ *74.5\ g/mol=208.6\ g" of KCl is dissolved per 1000 g of water.

"Solute/solvent\ mass\ ratio=\dfrac{208.6\ g}{1000\ g}=0.2086"


Let the mass of water in 300 grams of the solution be x, then the mass of KCl is 300-x and "\dfrac{300-x}{x}=0.2086;300-x=0.2086x;1.2086x=300;x=\dfrac{300}{1.2086}=248.22."